# Excess cataloged X-ray and radio sources at galaxy-cluster virial shocks

### 26 Feb 2024 v1

@@ 3.3 Poisson and control sample simulations
In the absence of any signals and source correlations, the statistical distribution of the small number of source counts in each radial bin within each cluster would follow a Poisson distribution, with a source density fixed by the foreground. However, the statistical significance and uncertainty estimates in Eqs. (5)–(6), (10), (3.1)– (17), (19), and (23) invoke, for simplicity, normally-distributed foreground counts. We run Monte-Carlo Poisson simulations to test and correct for this non-Poissonian approximation. Such simulations do not, however, capture structures and correlations in the actual catalogs. We therefore study, in addition, the statistics of the actual catalog sources around a large number of control (or mock) cluster samples, in order to test for systematics and validate our method.
@@
As expected, the results approach Eqs. (13) and (17) at large $\tau$, where $\mathcal{F}\gg 1$, but deviate substantially at small $\tau$. The deviations are particularly large for $\mathcal{N}<\mathcal{F}$ source deficits, but we are mostly interested in the $S>0$ significance of a positive excess. The approximate agreement of the $S=0,+1,+2\ldots$ curves with the corresponding Poisson and control-sample containment fractions confirms that our nominal §3.1 and §3.2 estimates capture the significance of the excess reasonably well, with small errors $\Delta S$ in $S$. Quantitatively, for the relevant, $\tau>1$ range, we find modest, $|\Delta S|/S\lesssim 10\%$ fractional errors, except for radio CW stacking where $|\Delta S|/S\lesssim 15\%$. While small, this $\Delta S$ should be accounted for at the VS region.
Much of the $\Delta S$ difference is attributed to the inaccurate normal-distribution approximation of Poisson statistics. Indeed, the Poisson and control-sample estimates are in much better agreement with each other, for both positive and negative $S$, and even for $\tau<1$. The difference $\Delta S$ between them is very small in X-rays, where $|\Delta S|/S\simeq 4\%$ ($|\Delta S|/S\simeq 2\%$) for $S>0$ ($S<0$), and is acceptable for radio SW stacking, where $|\Delta S|/S\simeq 6\%$ ($|\Delta S|/S\simeq 2\%$). For these cases, we henceforth apply an analytic correction (§C), correcting our nominal-distribution estimates for Poisson statistics; the relevant Eqs. (35)–(36) are shown to accurately follow the simulated Poisson results in Fig. 3. The overall agreement between Poisson and control-sample results indicates that systematic effects and correlations are weak, and our Poisson-statistics estimates are reliable.
@@ 3.4 Stacked source-excess models
$n_{\rm{planar}}\left(\tau\right)\equiv\frac{n_{0}\tau_{\rm sh}}{\sqrt{8\pi\,\Delta\tau_{\rm sh}^{2}}}\exp\left[-\frac{\left(\tau-{\tau_{\rm sh}}\right)^{2}}{2\,\Delta\tau_{\rm sh}^{2}}\right]\mbox{ .}$ (24)
@@
$n_{\rm shell}\left(\tau\right)\equiv\frac{n_{0}\tau_{\rm sh}\Theta\left(\tau_{\rm sh}-\tau\right)}{2\sqrt{\tau_{\rm sh}^{2}-\tau^{2}}}\,,$ (26)
where $\Theta$ is the Heaviside step function. The normalization is chosen to again satisfy Eq. (25), to facilitate a meaningful comparison of $n_{0}$ in the two models.
@@ 3.5 Model fitting and TS-Test
$\chi^{2}\simeq\sum_{\tau,{\mathsf{c}}}\frac{\left[\mathcal{N}(\tau,{\mathsf{c}})-\mathcal{F}(\tau,{\mathsf{c}})-\mathcal{M}(\tau,{\mathsf{c}})\right]^{2}}{\mathcal{F}(\tau,{\mathsf{c}})+\mathcal{M}(\tau,{\mathsf{c}})}\mbox{ ,}$ (27)
@@
$\chi^{2}_{\rm\mbox{\tiny{SW}}}\simeq\sum_{\tau}\frac{\left[\mathcal{N}(\tau)-\mathcal{F}(\tau)-\mathcal{M}(\tau)\right]^{2}}{\mathcal{F}(\tau)+\mathcal{M}(\tau)}\mbox{ ,}$ (28)
@@
$\chi^{2}_{\rm\mbox{\tiny{CW}}}$ $\simeq\sum_{\tau}\frac{\left[\sum\limits_{{\mathsf{c}}=1}^{{N}_{c}}\frac{\mathcal{N}(\tau,{\mathsf{c}})-\mathcal{F}(\tau,{\mathsf{c}})-\mathcal{M}(\tau,{\mathsf{c}})}{\sqrt{\mathcal{F}(\tau,{\mathsf{c}})}}\right]^{2}}{\sum\limits_{{\mathsf{c}}=1}^{{N}_{c}}\left[1+\frac{\mathcal{M}(\tau,{\mathsf{c}})}{\mathcal{F}(\tau,{\mathsf{c}})}\right]}\mbox{ .}$ (29)
@@
$\mbox{TS}\equiv-2\ln\frac{\mathcal{L}_{\rm max,-}}{\mathcal{L}_{\rm max,+}}=\chi^{2}_{-}-\chi^{2}_{+}\,,$ (30)
@@
For modelling, we nominally adopt a $\Delta\tau=1/6$ resolution, to better capture the radial profile; results for other resolutions are given in §D. We use Monte-Carlo simulations to verify that although the data is not normally distributed, $\chi^{2}$ in both Eqs. (28) and (29) indeed follows the expected $\chi^{2}$ distribution (see §C), and so can be used to estimate the TS-based CLs, denoted ${\sigma_{\tiny{\mbox{TS}}}}$.
@@ 4.2 Modeling
  • Nominal (SW; $\Delta\tau=1/6$) best-fitted parameters; additional models are provided in Table 3. Table columns: (1) Effective VS model (planar or shell); (2) Stacked catalog (2RXS, NVSS, or the two catalogs jointly); (3) and (4) Normalized, $n_{0}^{*}\equiv 10^{3}n_{0}$ X-ray and radio source densities in Eqs. (24) or (26); (5) Dimensionless shock radius; (6) Dimensionless signal width; (7) TS value; (8) TS-based significance.
@@ 5 Properties of VS excess sources
The $\delta\simeq 0.5$ fractional overdensity of both X-ray and radio catalog sources within the narrow VS range indicates that roughly $\sim 1/3$ of the sources therein are associated with the VS, while the remaining $\sim 2/3$ majority are field sources. Hence, while we cannot attribute individual sources to the VS, the most likely properties of those VS-induced sources can be deduced from a differential statistical analysis that compares the properties of sources in the VS region to their field counterparts.
@@
When relevant, we show analogous X-ray (left panels) and radio (right panels) ${\mathbb{P}}$ properties on the same figure. Each figure presents both differential $\mathbb{N}({\mathbb{P}})\equiv dN/d{\mathbb{P}}$ (top panels) and cumulative $N(<{\mathbb{P}})$ (bottom panels) distributions in the VS region (solid green), compared to equivalent field (dashed and dotted red) and ICM distributions scaled to the same $\Omega$. Bottom panels highlight the cumulative, $N_{\mbox{\tiny VS}}(<{\mathbb{P}})-N_{\mbox{\tiny F}}(<{\mathbb{P}})$ distribution (dashed blue) and $S({\mathbb{P}})=[\mathbb{N}_{\mbox{\tiny VS}}-\mu(\mathbb{N}_{\mbox{\tiny F}})]/\sigma(\mathbb{N}_{\mbox{\tiny F}})$ local significance (solid green) of the VS excess, with respect to the field distribution $\mathbb{N}_{\mbox{\tiny F}}$ inferred from $>200$ random field samples.
@@
Assigning sources to the redshift of the nearby MCXC cluster, as seen in projection, produces an intrinsic, albeit noisy, estimate of the typical luminosity of excess VS sources. Figure 8 thus shows that the excess is associated with $10^{42}\mbox{ erg}\mbox{ s}^{-1}\lesssim L_{X}\lesssim 10^{43}\mbox{ erg}\mbox{ s}^{-1}$ 2RXS sources and with $10^{40}\mbox{ erg}\mbox{ s}^{-1}\lesssim\nu L_{\nu}\lesssim 10^{41}\mbox{ erg}\mbox{ s}^{-1}$ NVSS sources. (The luminosity of excess GMRT sources is a factor $\sim 2$ lower than in NVSS; see Fig. E.6.) Again, local excess signals at lower or higher luminosities cannot be substantiated, as they are associated with few sources and are not robust against changes to the analysis details.
Figure 7: Differential $\mathbb{N}({\mathbb{P}})$ (top panels) and cumulative $N(<{\mathbb{P}})$ (bottom panels with left axes) distributions of logarithmic, ${\mathbb{P}}=\log_{10}F$ energy flux in X-rays (left panels) and radio (right panels). Differential distributions are shown for the VS region (solid green) and comparable ICM (dot-dashed magenta) and field (red; thick dashed for mean value and thin dotted for $1\sigma$ dispersion) estimates; see text. Bottom panels show, in addition to the cumulative distributions in the VS region (dot-dashed green), field estimate (dotted red), and their difference (dashed thick blue), also the local significance of the VS excess (solid green with right axes). In radio, we add a $\sigma_{\rm smooth}^{2}\equiv(0.2\,F_{\nu})^{2}$ smoothing variance to the $F_{\nu}$ uncertainty, for visibility.
@@
$\alpha\equiv-\frac{d\ln L}{d\ln\nu}\simeq 0.81\left(\frac{100\nu L_{\nu}}{L_{X}}\right)^{0.067}\,,$ (31)
providing a robust reference value for a spectral analysis.
@@
This estimate is broadly consistent but somewhat harder than in Eq. (31), and in the absence of NVSS-2RXS pairs (see §F), we cannot support such $\alpha_{\nu}$ values persisting out to X-rays. Moreover, while we cannot localize any excess in the X-ray spectral index $\Gamma_{X}$ obtained from power-law fits, both Mekal and black-body fits suggest $k_{B}T\gtrsim 1\mbox{ keV}$ excess-source temperatures, as shown in Figs. 10 and A.2, where $k_{B}$ is the Boltzmann constant.
Figure 10: Distribution of the ${\mathbb{P}}=\log_{10}k_{B}T$ Mekal-fitted temperature; notations as in Fig. 7.
@@ 6 Origin of the VS sources
The $\Delta P\lesssim 0.02\mbox{ eV}\mbox{ cm}^{-3}$ pressure change induced by the VS [e.g., ArnaudEtAl10, PlanckV13, HurierEtAl19, keshet20coincident] is much too weak to penetrate down to the vicinity of compact bodies, certainly those found in the centers of galaxies. As anticipated in §1, our results thus challenge previous claims that a well-localized excess of X-ray sources near the virial radius could be attributed to AGN. Instead, sources triggered by the VS are likely to be extended and gaseous. Indeed, this conclusion is directly corroborated by the inferred, 10101010100100100100 kpc extent of the excess sources, as well as by the polarization angle of the radio emission.
@@
$l_{\mbox{\tiny cool}}\simeq v_{s}t_{\mbox{\tiny cool}}\simeq v_{s}\frac{6\pi m_{e}cf_{B}}{\sigma_{T}\gamma_{\nu}B^{2}}\simeq 110\frac{\sqrt{B_{1}T_{1}}}{1+0.1B_{1}^{2}}\mbox{ kpc}$ (32)
@@
$f_{B}\equiv\frac{B^{2}/8\pi}{u_{cmb}+B^{2}/8\pi}\simeq\frac{0.1B_{1}^{2}}{1+0.1B_{1}^{2}}$ (33)
@@
In contrast to radio, the $\epsilon=\epsilon_{1}\mbox{ keV}$ X-ray sources are unlikely to arise from non-thermal, neither IC nor synchrotron, emission. Indeed, IC emission over the relevant scales, dominated by $\gamma_{\epsilon}\simeq 10^{3}\epsilon_{1}^{1/2}$ CREs scattering CMB photons, would extend over a far too long, $l_{\mbox{\tiny cool}}\simeq 2(T_{1}/\epsilon_{1})^{1/2}\mbox{ Mpc}$ cooling length, greatly exceeding the width of the signal; IC emission is also too faint to explain the sources, as shown below. The opposite problem of $l_{\mbox{\tiny cool}}\lesssim 9(T_{1}B_{1}/\epsilon_{1})^{1/2}\mbox{ pc}$ cooling and $l_{\mbox{\tiny acc}}<cl_{\mbox{\tiny cool}}/v_{s}$ acceleration lengths much too short to explain the signal would arise if X-rays were synchrotron emission, also requiring a radio spectrum similar to Eq. (31), which is softer than inferred in §5, as well as an ongoing acceleration of excessively energetic, $\gamma\gtrsim 10^{8.5}(\epsilon_{1}/B_{1})^{1/2}$ CREs. In either case, if CREs were responsible for the X-ray sources, we would expect to find at least a few 2RXS–NVSS pairs, absent in the present data.

### 19 Sep 2024 v2

@@ 3.3 Poisson and control sample simulations
In the absence of any signals and source correlations, the statistical distribution of the small number of source counts in each radial bin within each cluster would follow a Poisson distribution, with a source density fixed by the foreground. However, the statistical significance and uncertainty estimates in Eqs. (3.5)–(3.6), (3.10), (3.1)– (3.17), (3.19), and (3.23) invoke, for simplicity, normally-distributed foreground counts. We run Monte-Carlo Poisson simulations to test and correct for this non-Poissonian approximation. Such simulations do not, however, capture structures and correlations in the actual catalogs. We therefore study, in addition, the statistics of the actual catalog sources around a large number of control (or mock) cluster samples, in order to test for systematics and validate our method.
@@
As expected, the results approach Eqs. (3.13) and (3.17) at large $\tau$, where $\mathcal{F}\gg 1$, but deviate substantially at small $\tau$. The deviations are particularly large for $\mathcal{N}<\mathcal{F}$ source deficits, but we are mostly interested in the $S>0$ significance of a positive excess. The approximate agreement of the $S=0,+1,+2\ldots$ curves with the corresponding Poisson and control-sample containment fractions confirms that our nominal §3.1 and §3.2 estimates capture the significance of the excess reasonably well, with small errors $\Delta S$ in $S$. Quantitatively, for the relevant, $\tau>1$ range, we find modest, $|\Delta S|/S\lesssim 10\%$ fractional errors, except for radio CW stacking where $|\Delta S|/S\lesssim 15\%$. While small, this $\Delta S$ should be accounted for at the VS region.
Much of the $\Delta S$ difference is attributed to the inaccurate normal-distribution approximation of Poisson statistics. Indeed, the Poisson and control-sample estimates are in much better agreement with each other, for both positive and negative $S$, and even for $\tau<1$. The difference $\Delta S$ between them is very small in X-rays, where $|\Delta S|/S\simeq 4\%$ ($|\Delta S|/S\simeq 2\%$) for $S>0$ ($S<0$), and is acceptable for radio SW stacking, where $|\Delta S|/S\simeq 6\%$ ($|\Delta S|/S\simeq 2\%$). For these cases, we henceforth apply an analytic correction (§C), correcting our nominal-distribution estimates for Poisson statistics; the relevant Eqs. (C.1)–(C.2) are shown to accurately follow the simulated Poisson results in Fig. 3. The overall agreement between Poisson and control-sample results indicates that systematic effects and correlations are weak, and our Poisson-statistics estimates are reliable.
@@ 3.4 Stacked source-excess models
$n_{\rm{planar}}\left(\tau\right)\equiv\frac{n_{0}\tau_{\rm sh}}{\sqrt{8\pi\,\Delta\tau_{\rm sh}^{2}}}\exp\left[-\frac{\left(\tau-{\tau_{\rm sh}}\right)^{2}}{2\,\Delta\tau_{\rm sh}^{2}}\right]\mbox{ .}$ (3.24)
@@
$n_{\rm shell}\left(\tau\right)\equiv\frac{n_{0}\tau_{\rm sh}\Theta\left(\tau_{\rm sh}-\tau\right)}{2\sqrt{\tau_{\rm sh}^{2}-\tau^{2}}}\,,$ (3.26)
where $\Theta$ is the Heaviside step function. The normalization is chosen to again satisfy Eq. (3.25), to facilitate a meaningful comparison of $n_{0}$ in the two models.
@@ 3.5 Model fitting and TS-Test
$\chi^{2}\simeq\sum_{\tau,{\mathsf{c}}}\frac{\left[\mathcal{N}(\tau,{\mathsf{c}})-\mathcal{F}(\tau,{\mathsf{c}})-\mathcal{M}(\tau,{\mathsf{c}})\right]^{2}}{\mathcal{F}(\tau,{\mathsf{c}})+\mathcal{M}(\tau,{\mathsf{c}})}\mbox{ ,}$ (3.27)
@@
$\chi^{2}_{\rm\mbox{\tiny{SW}}}\simeq\sum_{\tau}\frac{\left[\mathcal{N}(\tau)-\mathcal{F}(\tau)-\mathcal{M}(\tau)\right]^{2}}{\mathcal{F}(\tau)+\mathcal{M}(\tau)}\mbox{ ,}$ (3.28)
@@
$\chi^{2}_{\rm\mbox{\tiny{CW}}}$ $\simeq\sum_{\tau}\frac{\left[\sum\limits_{{\mathsf{c}}=1}^{{N}_{c}}\frac{\mathcal{N}(\tau,{\mathsf{c}})-\mathcal{F}(\tau,{\mathsf{c}})-\mathcal{M}(\tau,{\mathsf{c}})}{\sqrt{\mathcal{F}(\tau,{\mathsf{c}})}}\right]^{2}}{\sum\limits_{{\mathsf{c}}=1}^{{N}_{c}}\left[1+\frac{\mathcal{M}(\tau,{\mathsf{c}})}{\mathcal{F}(\tau,{\mathsf{c}})}\right]}\mbox{ .}$ (3.29)
@@
$\mbox{TS}\equiv-2\ln\frac{\mathcal{L}_{\rm max,-}}{\mathcal{L}_{\rm max,+}}=\chi^{2}_{-}-\chi^{2}_{+}\,,$ (3.30)
@@
For modelling, we nominally adopt a $\Delta\tau=1/6$ resolution, to better capture the radial profile; results for other resolutions are given in §D. We use Monte-Carlo simulations to verify that although the data is not normally distributed, $\chi^{2}$ in both Eqs. (3.28) and (3.29) indeed follows the expected $\chi^{2}$ distribution (see §C), and so can be used to estimate the TS-based CLs, denoted ${\sigma_{\tiny{\mbox{TS}}}}$.
@@ 4.2 Modeling
  • Nominal (SW; $\Delta\tau=1/6$) best-fitted parameters; additional models are provided in Table 3. Table columns: (1) Effective VS model (planar or shell); (2) Stacked catalog (2RXS, NVSS, or the two catalogs jointly); (3) and (4) Normalized, $n_{0}^{*}\equiv 10^{3}n_{0}$ X-ray and radio source densities in Eqs. (3.24) or (3.26); (5) Dimensionless shock radius; (6) Dimensionless signal width; (7) TS value; (8) TS-based significance.
@@ 5 Properties of VS excess sources
The $\delta\simeq 0.5$ fractional overdensity of both X-ray and radio catalog sources within the narrow VS range indicates that roughly $\sim 1/3$ of the sources therein are associated with the VS, while the remaining $\sim 2/3$ majority are field sources. The sources found in the VS regions, listed in §H, are mostly unidentified; note that these peripheral sources lie well outside the radii of known ICM radio sources, including halos, relics, and even mega-halos [Keshet24], and cannot be attributed to the smooth distribution of cluster sources or to AGN (§G). While we cannot attribute individual sources to the VS, the most likely properties of those VS-induced sources can be deduced from a differential statistical analysis that compares the properties of sources in the VS region to their field counterparts.
@@
When relevant, we show analogous X-ray (left panels) and radio (right panels) ${\mathbb{P}}$ properties on the same figure. Each figure presents both differential $\mathbb{N}({\mathbb{P}})\equiv dN/d{\mathbb{P}}$ (top panels) and cumulative $N(<{\mathbb{P}})$ (bottom panels) distributions in the VS region (solid and dot-dashed green), compared to equivalent field (thick dashed and thin dotted red) and ICM (dot-dashed and double dot-dashed magenta) distributions scaled to the same $\Omega$. Bottom panels highlight also the cumulative, $N_{\mbox{\tiny VS}}(<{\mathbb{P}})-N_{\mbox{\tiny F}}(<{\mathbb{P}})$ distribution (solid blue) and the $S({\mathbb{P}})=[\mathbb{N}_{\mbox{\tiny VS}}-\mu(\mathbb{N}_{\mbox{\tiny F}})]/\sigma(\mathbb{N}_{\mbox{\tiny F}})$ local significance (dashed green) of the VS excess, with respect to the field distribution $\mathbb{N}_{\mbox{\tiny F}}$ inferred from $>200$ random field samples.
@@
The cataloged sources do not have readily available associated redshifts. Assigning sources to the redshift of the nearby MCXC cluster, as seen in projection, produces an intrinsic, albeit noisy, estimate of the typical luminosity of excess VS sources. Figure 8 thus shows that the excess is associated with $10^{42}\mbox{ erg}\mbox{ s}^{-1}\lesssim L_{X}\lesssim 10^{43}\mbox{ erg}\mbox{ s}^{-1}$ 2RXS sources and with $10^{40}\mbox{ erg}\mbox{ s}^{-1}\lesssim\nu L_{\nu}\lesssim 10^{41}\mbox{ erg}\mbox{ s}^{-1}$ NVSS sources. (The luminosity of excess GMRT sources is a factor $\sim 2$ lower than in NVSS; see Fig. E.6.) Again, local excess signals at lower or higher luminosities cannot be substantiated, as they are associated with few sources and are not robust against changes to the analysis details.
$\alpha\equiv-\frac{d\ln L}{d\ln\nu}\simeq 0.81\left(\frac{100\nu L_{\nu}}{L_{X}}\right)^{0.067}\,,$ (5.1)
providing a robust reference value for a spectral analysis.
Figure 7: Differential $\mathbb{N}({\mathbb{P}})$ (top panels) and cumulative $N(<{\mathbb{P}})$ (bottom panels with left axes) distributions of logarithmic, ${\mathbb{P}}=\log_{10}F$ energy flux in X-rays (left panels) and radio (right panels). Differential distributions are shown for the VS region (solid green) and comparable ICM (dot-dashed magenta) and field (red; thick dashed for mean value and thin dotted for $1\sigma$ dispersion) estimates; see text. Bottom panels show, in addition to the cumulative distributions in the VS region (dot-dashed green), field estimate (dotted red), and their difference (solid blue), also the local significance of the VS excess (dashed green with right axes). In radio, we add a $\sigma_{\rm smooth}^{2}\equiv(0.2\,F_{\nu})^{2}$ smoothing variance to the $F_{\nu}$ uncertainty, for visibility.
@@
This estimate is broadly consistent but somewhat harder than in Eq. (5.1), and in the absence of NVSS-2RXS pairs (see §F), we cannot support such $\alpha_{\nu}$ values persisting out to X-rays. Moreover, while we cannot localize any excess in the X-ray spectral index $\Gamma_{X}$ obtained from power-law fits, both Mekal and black-body fits suggest $k_{B}T\gtrsim 1\mbox{ keV}$ excess-source temperatures, as shown in Figs. 10 and A.2, where $k_{B}$ is the Boltzmann constant. The combined implications of hard power-law spectra in radio and thermal spectra in X-rays are discussed in §6.
@@
Figure 10: Distribution of the ${\mathbb{P}}=\log_{10}k_{B}T$ Mekal-fitted temperature; notations as in Fig. 7. Here we show also the cumulative ICM-source distribution (double-dot magenta).
@@
This effect is quite robust, strengthening at high Galactic latitudes or at high declinations [where the typical Galactic Faraday rotation at NVSS frequencies drops below $\sim 30^{\circ}$; see, e.g., TaylorEtAl09], as illustrated by the figure. Such a polarization supports the association of the sources not only with the cluster, but also specifically with its VS, suggesting that the emission arises from a gaseous medium shocked by the VS, resulting in enhanced tangential magnetic fields. While the measured polarizations carry substantial uncertainties, the elevated polarized flux and the correlation between polarization vectors and radii among VS (but not field) sources suggest that the polarization effect is real.
@@ 6 Origin of the VS sources
The $\Delta P\lesssim 0.02\mbox{ eV}\mbox{ cm}^{-3}$ pressure change induced by the VS [e.g., ArnaudEtAl10, PlanckV13, HurierEtAl19, keshet20coincident] is much too weak to penetrate down to the vicinity of compact bodies, certainly those found in the centers of galaxies. As anticipated in §1, our results thus challenge previous claims that a well-localized excess of X-ray sources near the virial radius could be attributed to AGN. Indeed, sources already associated with identified AGN show no significant localized excess; see §G. Instead, sources triggered by the VS are likely to be extended and gaseous. Indeed, this conclusion is directly corroborated by the inferred, 10101010100100100100 kpc extent of the excess sources, as well as by the polarization angle of the radio emission.
@@
$l_{\mbox{\tiny cool}}\simeq v_{s}t_{\mbox{\tiny cool}}\simeq v_{s}\frac{6\pi m_{e}cf_{B}}{\sigma_{T}\gamma_{\nu}B^{2}}\simeq 110\frac{\sqrt{B_{1}T_{1}}}{1+0.1B_{1}^{2}}\mbox{ kpc}$ (6.1)
@@
$f_{B}\equiv\frac{B^{2}/8\pi}{u_{cmb}+B^{2}/8\pi}\simeq\frac{0.1B_{1}^{2}}{1+0.1B_{1}^{2}}$ (6.2)
@@
In contrast to radio, the $\epsilon=\epsilon_{1}\mbox{ keV}$ X-ray sources are unlikely to arise from non-thermal, neither IC nor synchrotron, emission. Indeed, IC emission over the relevant scales, dominated by $\gamma_{\epsilon}\simeq 10^{3}\epsilon_{1}^{1/2}$ CREs scattering CMB photons, would extend over a far too long, $l_{\mbox{\tiny cool}}\simeq 2(T_{1}/\epsilon_{1})^{1/2}\mbox{ Mpc}$ cooling length, greatly exceeding the width of the signal; IC emission is also too faint to explain the sources, as shown below. The opposite problem of $l_{\mbox{\tiny cool}}\lesssim 9(T_{1}B_{1}/\epsilon_{1})^{1/2}\mbox{ pc}$ cooling and $l_{\mbox{\tiny acc}}<cl_{\mbox{\tiny cool}}/v_{s}$ acceleration lengths much too short to explain the signal would arise if X-rays were synchrotron emission, also requiring a radio spectrum similar to Eq. (5.1), which is softer than inferred in §5, as well as an ongoing acceleration of excessively energetic, $\gamma\gtrsim 10^{8.5}(\epsilon_{1}/B_{1})^{1/2}$ CREs. In either case, if CREs were responsible for the X-ray sources, we would expect to find at least a few 2RXS–NVSS pairs, absent in the present data.